# bibliography.bib

### v2

@@ 4.2 Representations of $\operatorname{GL}(m+nP)$
By Corollary 3.13, $G=\operatorname{GL}(m+nP)$ has a Gauss decomposition $(\mathbb{G}^{\prime}_{a})^{mn}\times B_{m,n}$. For ease of notation, let $B:=B_{m,n}$ and $M:=(\mathbb{G}^{\prime}_{a})^{mn}$. Thus, every irreducible $G$-module has a “highest weight”, which is the corresponding irreducible $B$-module, and this map from $G$-irreps to $B$-irreps is injective. For any irreducible $B$-module $V$, we have $\operatorname{Hom}_{B}(V,\mathcal{O}(G))\neq 0$ iff there exists a $G$-module with highest weight $V$. The Gauss decomposition implies that the multiplication map is an isomorphism $\mathcal{O}(M)\otimes\mathcal{O}(B)\cong\mathcal{O}(G)$. Therefore,
@@
and every irreducible $B$-module is the highest weight for some irreducible $G$-module. We thus get a bijection between irreducible $G$-modules and irreducible $B$-modules, which are in bijection with pairs of finite-dimensional irreducible $\operatorname{GL}(m)$ and $\operatorname{GL}(nP)$ modules. ∎
Thus, to describe irreducible $\operatorname{GL}(m+nP)$-modules, it suffices to describe irreducible $\operatorname{GL}(nP)$ modules. We first find all finite-dimensional irreducible $\operatorname{GL}(P)$-modules (Subsection ), and use these to describe the irreducible $\operatorname{GL}(nP)$-modules in Subsection 4.2.3.
4.2.1 Representations of $\operatorname{GL}(P)$
@@
On $P$, $(a,b)$ takes $x\mapsto x\otimes a+x^{\prime}\otimes b$. So
$x^{n}\mapsto(x\otimes a+x^{\prime}\otimes b)^{n}=(x\otimes a)^{n}+\sum_{k=0}^{n-1}(x\otimes a)^{k}\otimes(x^{\prime}\otimes b)\otimes(x\otimes a)^{n-k-1}.$
@@
$(x\otimes a+x^{\prime}\otimes b)^{n}=x^{n}\otimes a^{n}+x^{n-1}x^{\prime}\otimes\left(nba^{n-1}+\frac{n(n-1)}{2}a^{n-2}a^{\prime}(a+b^{\prime})\right).$ (3)
@@
We can now describe the Ext quiver for the degree 0 block.
@@
Corollary 4.13.
$\operatorname{Ext}(\mathds{1},\mathds{1})\cong\operatorname{Ext}(\xi,\xi)$ and $\operatorname{Ext}(\mathds{1},\xi)\cong\operatorname{Ext}(\xi,\mathds{1})$ via tensoring an extension by $\xi$.
Hence it suffices to compute primitive elements and skew-primitive elements $f\in\mathcal{O}(\operatorname{GL}(P))$ such that $\Delta(f)=1\otimes f+f\otimes\xi$.
@@
$\begin{pmatrix}1&0\\ ba^{-1}&1+(ba^{-1})^{\prime}\end{pmatrix}\mapsto\begin{pmatrix}1&\lambda_{1}(ba^{-1})^{\prime}+\lambda_{2}ba^{-1}\\ 0&1+(ba^{-1})^{\prime}\end{pmatrix}$
@@
In addition to this explicit computation of the representations, we can see how they restrict to representations of $M_{1}$ and $H_{1}$. Recall that $\operatorname{GL}(P)=M_{1}\times H_{1}$ via
$\begin{pmatrix}a&a^{\prime}\\ b&a+b^{\prime}\end{pmatrix}=\begin{pmatrix}1&0\\ ba^{-1}&1+(ba^{-1})^{\prime}\end{pmatrix}\begin{pmatrix}a&a^{\prime}\\ 0&a\end{pmatrix}.$
@@
Recall that $L_{\xi}$ is the irreducible $M_{1}$-representation defined by $a\mapsto 1+a^{\prime}$. The invariant functions we get in $\operatorname{Hom}_{M_{1}}(L_{\xi},\mathcal{O}(G))$ generate the $G$-representation $\xi$; indeed, $\xi(a,b)=1+(ba^{-1})^{\prime}$. Recall that $L_{n}$ is the irreducible $H_{1}$-representation corresponding to the intgere $n$. We can check that the $H_{1}$-invariant functions $L_{n}\to\mathcal{O}(G)$ corresponding to $L_{n}^{*}\otimes 1$ generate either $T_{n}$ or $\chi^{n/4}$ depending on whether $n\equiv 0\pmod{4}$ or not.
4.2.2 Explicit structure of representations of $\operatorname{GL}(\mathds{1}+P)$ and their tensor products
@@
Table 1: Weight decompositions for tensor products of irreducible representation of $\operatorname{GL}(1+P)$
4.2.3 Representations of $\operatorname{GL}(nP)$
@@
where $X_{ij}\in\operatorname{End}_{A}(P\otimes A)$, we see that $\operatorname{GL}(nP)$ has a subgroup $B$ consisting of the upper triangular block matrices of this form with the $X_{ii}$ invertible. The irreducible representations of this subgroup $B$ are $n$-tuples of irreducible $\operatorname{GL}(P)$-modules, and the irreducible representations of $\operatorname{GL}(nP)$ correspond to the irreducible $B$-modules $V$ with $\operatorname{Hom}_{B}(V,\mathcal{O}(\operatorname{GL}(nP)))\neq 0$.
Using the same degree argument as in the $\operatorname{GL}(n)$ case, we see that $(L_{1},\dots,L_{n})$ can only correspond to a $\operatorname{GL}(nP)$-module if $\deg L_{1}\geq\cdots\geq\deg L_{n}$.
@@
Let $G=\operatorname{GL}(nP)$ and let $N$ be the nilradical of $\mathcal{O}(G)$; then $N$ is the ideal generated by $\operatorname{im}d$ and $\mathcal{O}(G)/N=\mathcal{O}(G_{*})$ for $G_{*}$ the closed subgroup such that $G_{*}(A)=\left\{\begin{pmatrix}C&0\\ D&C\end{pmatrix}\right\}$, $C\in\operatorname{GL}(n)(A)$, $D\in\mathfrak{gl}_{n}(A)$, so $dC=dD=0$. Multiplication on $G_{*}$ is given by
@@
$\delta_{p}(I^{m+1})=0,\delta_{p}(p)=1,\delta_{p}(q)=1$
@@
Combining this with 4.5, we get our main result:
Theorem 4.21.
4.2.4 An alternative construction of representations of $\operatorname{GL}(nP)$
@@
Question 4.28.
What are the blocks of $\operatorname{Rep}(\operatorname{GL}(m+nP))$?
4.2.5 The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
While the Frobenius map $a\mapsto a^{2}$ is not a homomorphism in $\operatorname{Ver}_{4}^{+}$, it is true that $a\mapsto a^{4}$ is a homomorphism. In particular, this induces a homomorphism $\operatorname{Fr}^{2}:\operatorname{GL}(nP)\to\operatorname{GL}(n)$ which takes a pair of matrices $B,C\in\operatorname{Mat}_{n\times n}(A)$, $B$ invertible, to the matrix in $\operatorname{GL}(n)(A)$ with entries $b_{ij}^{4}$. The $\operatorname{GL}(n)$-irrep with highest weight $(\lambda_{1},\dots,\lambda_{n})$ pulls back to the $\operatorname{GL}(nP)$ irrep with highest weight $(\chi^{\lambda}_{1},\dots,\chi^{\lambda}_{n})$.
@@ 4.2.5 The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
Thus even though the Frobenius map itself does not have good properties, the square of the Frobenius map does, and we expect that a Steinberg tensor product theorem should hold for $\operatorname{GL}(nP)$ using the square of the Frobenius. Therefore, we propose the following conjecture. Say that a highest weight $\lambda$ of $\operatorname{GL}(nP)$ is $4$-restricted if $\deg\lambda_{i}-\deg\lambda_{i+1}<4$. Then every highest weight can be written as $\lambda\cdot\chi^{\mu}=(\chi^{\mu_{1}}\otimes\lambda_{1},\dots,\chi^{\mu_{n}}\otimes\lambda_{n})$ where $\lambda$ is $4$-restricted and $\mu=(\mu_{1},\dots,\mu_{n})$ is a highest $\operatorname{GL}(n)$ weight.
@@
For $\lambda$ 4-restricted and $\mu=(\mu_{1},\dots,\mu_{n})$ a highest $\operatorname{GL}(n)$-weight,
$L_{\operatorname{GL}(nP)}(\lambda\cdot\chi^{\mu})\cong L_{\operatorname{GL}(nP)}(\lambda)\otimes(\operatorname{Fr}^{2})^{*}L_{\operatorname{GL}(n)}(\mu).$

### v3

@@ 4.2. Representations of $\operatorname{GL}(m+nP)$
By Corollary 3.12, $G=\operatorname{GL}(m+nP)$ has a Gauss decomposition $(\mathbb{G}^{\prime}_{a})^{mn}\times B_{m,n}$. For ease of notation, let $B:=B_{m,n}$ and $M:=(\mathbb{G}^{\prime}_{a})^{mn}$. The irreducible $B$-modules are then in bijection with pairs of an irreducible $\operatorname{GL}(m)$-module and an irreducible $\operatorname{GL}(nP)$-module.
Via the discussion in Section 2.4, every irreducible $G$-module has a highest weight, which is the corresponding irreducible $B$-module, and this map from $G$-irreps to $B$-irreps is injective. For any irreducible $B$-module $V$, we have $\operatorname{Hom}_{B}(V,\mathcal{O}(G))\neq 0$ iff there exists a $G$-module with highest weight $V$. The Gauss decomposition implies that the multiplication map is an isomorphism $\mathcal{O}(M)\otimes\mathcal{O}(B)\cong\mathcal{O}(G)$. Therefore,
@@
and every irreducible $B$-module is the highest weight for some irreducible $G$-module. We thus get a bijection between irreducible $G$-modules and irreducible $B$-modules. ∎
Thus, to describe irreducible $\operatorname{GL}(m+nP)$-modules, it suffices to describe irreducible $\operatorname{GL}(nP)$ modules. We first find all finite-dimensional irreducible $\operatorname{GL}(P)$-modules (Section 4.2.1), and use these to describe the irreducible $\operatorname{GL}(nP)$-modules in Section 4.2.3. From that discussion, we will obtain the following theorem.
Theorem 4.6.
Question 4.7.
What are the blocks of $\operatorname{Rep}(\operatorname{GL}(m+nP))$?
4.2.1. Representations of $\operatorname{GL}(P)$
@@
Throughout this proof, we use the identification $\operatorname{Sym}^{n}(P)\otimes A=\operatorname{Sym}^{n}_{A}(P\otimes A)$ to explicitly describe the $\operatorname{GL}(P)$-action on $\operatorname{Sym}^{n}(P)$.
@@
On $P$, $(a,b)$ takes $x\otimes 1\mapsto x\otimes a+x^{\prime}\otimes b$. So
$x^{n}\otimes 1\mapsto(x\otimes a+x^{\prime}\otimes b)^{n}=(x\otimes a)^{n}+\sum_{k=0}^{n-1}(x\otimes a)^{k}\otimes(x^{\prime}\otimes b)\otimes(x\otimes a)^{n-k-1}.$
@@
(3) $(x\otimes a+x^{\prime}\otimes b)^{n}=x^{n}\otimes a^{n}+x^{n-1}x^{\prime}\otimes\left(nba^{n-1}+\frac{n(n-1)}{2}a^{n-2}a^{\prime}(a+b^{\prime})\right).$
@@
In particular, we can define a notion of degree on the irreducible $\operatorname{GL}(P)$-representations via the inclusion of $\mathbb{G}_{m}\subset\operatorname{GL}(P)$, $a\mapsto\begin{pmatrix}a&0\\ 0&a\end{pmatrix}$ for $a\in A^{\times}$: the degree of an irreducible $\operatorname{GL}(P)$-representation is its degree as a $\mathbb{G}_{m}$-representation. Hence, we can characterize $\operatorname{Rep}(\operatorname{GL}(P))$ as follows:
@@
We can now describe the Ext quiver for the degree 0 block. Let $G=\operatorname{GL}(P)$; then we have $\mathcal{O}(G)=k[A,B,A^{-1},A^{\prime},B^{\prime}]/(A^{\prime})^{2}=(B^{\prime})^{2}=0$, where $A$ corresponds to the function $(a,b)\mapsto a$ and $B$ to $(a,b)\mapsto b$. Then the characters of $G$ can be treated as elements of $\mathcal{O}(G)$, e.g. $\mathds{1}$ as $1\in\mathcal{O}(G)$ and $\xi$ as $1+(BA^{-1})^{\prime}\in\mathcal{O}(G)$.
@@
Since $\operatorname{Ext}(\mathds{1},\mathds{1})\cong\operatorname{Ext}(\xi,\xi)$ and $\operatorname{Ext}(\mathds{1},\xi)\cong\operatorname{Ext}(\xi,\mathds{1})$ via tensoring an extension by $\xi$, it suffices to compute primitive elements and skew-primitive elements $f\in\mathcal{O}(G)$ such that $\Delta(f)=1\otimes f+f\otimes\xi$.
@@
$\begin{pmatrix}1&0\\ ba^{-1}&1+(ba^{-1})^{\prime}\end{pmatrix}\mapsto\begin{pmatrix}1&\lambda_{1}(ba^{-1})^{\prime}+\lambda_{2}ba^{-1}\\ 0&1+(ba^{-1})^{\prime}\end{pmatrix}.$
@@
We can also see how these irreducible $\operatorname{GL}(P)$ representations restrict to representations of $M_{1}$ and $H_{1}$.
@@
Recall that $L_{\xi}$ is the irreducible $M_{1}$-representation defined by $a\mapsto 1+a^{\prime}$. The invariant functions we get in $\operatorname{Hom}_{M_{1}}(L_{\xi},\mathcal{O}(G))$ generate the $G$-representation $\xi$; indeed, $\xi(a,b)=1+(ba^{-1})^{\prime}$. Recall that $L_{n}$ is the irreducible $H_{1}$-representation corresponding to the integer $n$. We can check that the $H_{1}$-invariant functions $L_{n}\to\mathcal{O}(G)$ corresponding to $L_{n}^{*}\otimes 1$ generate either $T_{n}$ or $\chi^{n/4}$ depending on whether $n\equiv 0\pmod{4}$ or not.
4.2.2. Explicit structure of representations of $\operatorname{GL}(\mathds{1}+P)$ and their tensor products
@@
Table 1. Weight decompositions for tensor products of irreducible representation of $\operatorname{GL}(1+P)$
4.2.3. Representations of $\operatorname{GL}(nP)$
@@
where $X_{ij}\in\operatorname{End}_{A}(P\otimes A)$, we see that $\operatorname{GL}(nP)$ has a subgroup $B$ consisting of the upper triangular block matrices of this form with the $X_{ii}$ invertible. The irreducible representations of this subgroup $B$ are $n$-tuples of irreducible $\operatorname{GL}(P)$-modules, and per the discussion in Section 2.4, every irreducible representation of $\operatorname{GL}(nP)$ has a unique $B$-highest weight, which is an irreducible $B$-module $V$. The irreducible representations of $\operatorname{GL}(nP)$ correspond to the irreducible $B$-modules $V$ with $\operatorname{Hom}_{B}(V,\mathcal{O}(\operatorname{GL}(nP)))\neq 0$.
Lemma 4.19.
If the $n$-tuple of $\operatorname{GL}(P)$-irreps $(L_{1},\dots,L_{n})$ does not satisfy $\deg L_{1}\geq\cdots\geq\deg L_{n}$, then it cannot be the highest weight of an irreducible $\operatorname{GL}(nP)$-module.
Proof.
We have $\operatorname{GL}(n)\subset\operatorname{GL}(nP)$ via sending an invertible matrix $M\in\operatorname{GL}(n)(A)$ to $\begin{pmatrix}M&0\\ 0&M\end{pmatrix}$. A $\operatorname{GL}(nP)$-module with highest weight $(L_{1},\dots,L_{n})$ restricts to a $\operatorname{GL}(n)$-module (not necessarily irreducible) with highest weight $\deg L_{1},\dots,\deg L_{n}$, which cannot be finite-dimensional unless $\deg L_{1}\geq\cdots\geq\deg L_{n}$. ∎
@@
Let $G=\operatorname{GL}(nP)$ and let $N$ be the nilradical of $\mathcal{O}(G)$; then $N$ is the ideal generated by $\operatorname{im}d$ and $\mathcal{O}(G)/N=\mathcal{O}(G_{*})$ for $G_{*}$ the closed subgroup such that $G_{*}(A)=\left\{\begin{pmatrix}C&0\\ D&C\end{pmatrix}\right\}$, $C\in\operatorname{GL}(n)(A)$, $D\in\mathfrak{gl}_{n}(A)$, so $dC=dD=0$. Let $(D,C)\in G_{*}(A)$ denote the element
$\begin{pmatrix}I&0\\ D&I\end{pmatrix}\begin{pmatrix}C&0\\ 0&C\end{pmatrix}.$
Then multiplication on $G_{*}$ is given by
@@
$\delta_{p}(I^{m+1})=0,\delta_{p}(p)=1,\delta_{p}(q)=0$
@@
4.2.4. An alternative construction of representations of $\operatorname{GL}(nP)$
@@
4.2.5. The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
While the Frobenius map $a\mapsto a^{2}$ is not a homomorphism in $\operatorname{Ver}_{4}^{+}$, it is true that $a\mapsto a^{4}$ is a homomorphism. In particular, this induces a homomorphism $\operatorname{Fr}_{\text{odd}}:\operatorname{GL}(m+nP)\to\operatorname{GL}(n)$, acting on an element of $\operatorname{GL}(m+nP)(A)$ as
$\begin{pmatrix}F&C&C^{\prime}\\ B^{\prime}&D&D^{\prime}\\ B&E&D+E^{\prime}\end{pmatrix}\mapsto D^{(2)}$
where $D^{(2)}$ denotes the matrix in $\operatorname{GL}(n)(A)$ with entries $d_{ij}^{4}$.
We also have a homomorphism $\operatorname{Fr}_{\text{even}}:\operatorname{GL}(m+nP)\to\operatorname{GL}(m)$ which sends an element of $\operatorname{GL}(m+nP)(A)$ as above to $F^{(1)}$ (which is the matrix with entries $f_{ij}^{2}$).
Proposition 4.29.
$\operatorname{Fr}_{\text{even}}$ is a homomorphism.
Proof.
Let us compute the $ij$th matrix entry of
$\operatorname{Fr}_{\text{even}}\left(\begin{pmatrix}F_{1}&C_{1}&C_{1}^{\prime}\\ B_{1}^{\prime}&D_{1}&D_{1}^{\prime}\\ B_{1}&E_{1}&D_{1}+E_{1}^{\prime}\end{pmatrix}\begin{pmatrix}F_{2}&C_{2}&C_{2}^{\prime}\\ B_{2}^{\prime}&D_{2}&D_{2}^{\prime}\\ B_{2}&E_{2}&D_{2}+E_{2}^{\prime}\end{pmatrix}\right)=(F_{1}F_{2}+C_{1}B_{2}^{\prime}+C_{1}^{\prime}B_{2})^{(1)}.$
Denote the matrix entries of $F_{k}$ as $f^{k}_{ij}$ (all of which are in the kernel of $d$), the matrix entries of $C_{1}$ as $c_{ij}$, and the matrix entries of $B_{2}$ as $b_{ij}$. Then
$(F_{1}F_{2}+C_{1}B_{2}^{\prime}+C_{1}^{\prime}B_{2})^{(1)}_{ij}=\left(\sum_{k}f^{1}_{ik}f^{2}_{kj}+c_{ik}b_{kj}^{\prime}+c^{\prime}_{ik}b_{kj}\right)^{2}.$
We have that each $f^{1}_{ik}f^{2}_{kj}$, being in $\ker d$, commutes with all other terms, while each $c_{ik}b_{kj}^{\prime}$ and $c_{ik}^{\prime}b_{kj}$ squares to 0. Note also that
$c_{ik}b^{\prime}_{kj}c_{ik^{\prime}}b^{\prime}_{k^{\prime}j}+c_{ik^{\prime}}b^{\prime}_{k^{\prime}j}c_{ik}b^{\prime}_{kj}=c^{\prime}_{ik^{\prime}}c_{ik}^{\prime}b^{\prime}_{kj}b^{\prime}_{k^{\prime}j}=c^{\prime}_{ik}b_{kj}c^{\prime}_{ik^{\prime}}b_{k^{\prime}j}+c^{\prime}_{ik^{\prime}}b_{k^{\prime}j}+c^{\prime}_{ik}b_{kj}.$
Likewise,
$c_{ik}b^{\prime}_{kj}c^{\prime}_{ik^{\prime}}b_{k^{\prime}j}+c^{\prime}_{ik^{\prime}}b_{k^{\prime}j}c_{ik}b^{\prime}_{kj}=c^{\prime}_{ik}b_{kj}c_{ik^{\prime}}b^{\prime}_{k^{\prime}j}+c_{ik^{\prime}}b^{\prime}_{k^{\prime}j}c^{\prime}_{ik}b_{kj}.$
Hence, all terms except those in $(F_{1}F_{2})^{(1)}$ cancel, so
$(F_{1}F_{2}+C_{1}B_{2}^{\prime}+C_{1}^{\prime}B_{2})^{(1)}=(F_{1}F_{2})^{(1)}.$
Because both of these maps correspond to injective maps on their coordinate rings ($\operatorname{Fr}_{\text{odd}}$ sends matrix functions to their fourth powers, and $\operatorname{Fr}_{\text{even}}$ to their squares), they are surjections of schemes. Therefore, the pullback of an irreducible $\operatorname{GL}(m)$ or $\operatorname{GL}(n)$-module along one of these homomorphisms is an irreducible $\operatorname{GL}(m+nP)$-module. By checking highest weights, we have
$\operatorname{Fr}_{\text{even}}^{*}(L_{\operatorname{GL}(m)}(\lambda))$ $=L(2\lambda,\mathds{1})$
$\operatorname{Fr}_{\text{odd}}^{*}(L_{\operatorname{GL}(n)}(\lambda))$ $=L(0,\chi^{\mu})$
where $\chi^{\mu}=(\chi^{\mu_{1}},\dots,\chi^{\mu_{n}})$.
Remark 4.30.
The terminology of an even and odd Frobenius comes from the following: for an object $X=V\oplus W\otimes P\in\operatorname{Ver}_{4}^{+}$, notice that $\operatorname{Fr}_{\text{odd}}$ corresponds to the functor (which is not a tensor functor) $X\mapsto W^{(2)}$, where $W^{(2)}$ is the second Frobenius twist of $W$, while $\operatorname{Fr}_{\text{even}}$ corresponds to the functor $X\mapsto V^{(1)}=X^{(1)}$, the usual Frobenius twist of $X$.
@@ 4.2.5. The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
We expect that a Steinberg tensor product theorem should hold for $\operatorname{GL}(m+nP)$ using both $\operatorname{Fr}_{\text{odd}}$ and $\operatorname{Fr}_{\text{even}}$. Say that a highest weight $\Lambda$ of $\operatorname{GL}(nP)$ is $4$-restricted if $\deg\Lambda_{i}-\deg\Lambda_{i+1}<4$. Then every highest weight of $\operatorname{GL}(nP)$ can be written as $\Lambda\cdot\chi^{\mu}=(\chi^{\mu_{1}}\otimes\Lambda_{1},\dots,\chi^{\mu_{n}}\otimes\Lambda_{n})$ where $\Lambda$ is $4$-restricted and $\mu=(\mu_{1},\dots,\mu_{n})$ is a highest $\operatorname{GL}(n)$ weight. Say that a highest weight $(\mu,\Lambda)$ of $\operatorname{GL}(m+nP)$ is restricted if $\mu$ is 2-restricted and $\Lambda$ is 4-restricted as a $\operatorname{GL}(nP)$-highest weight. Our conjecture states that every irreducible representation of $\operatorname{GL}(m+nP)$ can be written as a tensor product of $L(\mu,\Lambda)$ for $(\mu,\Lambda)$ restricted with Frobenius pullbacks of irreducible $\operatorname{GL}(m)$ and $\operatorname{GL}(n)$ modules.
@@
For $\lambda$ a 4-restricted $\operatorname{GL}(nP)$ weight, $\mu$ a 2-restricted $\operatorname{GL}(m)$-weight, $\nu$ a $\operatorname{GL}(n)$-weight, and $\rho$ a $\operatorname{GL}(m)$-weight, we have
$L_{\operatorname{GL}(m+nP)}(\mu+2\rho,\lambda\cdot\chi^{\nu})\cong L_{\operatorname{GL}(m+nP)}(\mu,\lambda)\otimes(\operatorname{Fr}_{\text{odd}})^{*}L_{\operatorname{GL}(n)}(\nu)\otimes(\operatorname{Fr}_{\text{even}})^{*}L_{\operatorname{GL}(m)}(\rho).$